# Uncovering LLM-Generated Code: A Zero-Shot Synthetic Code Detector via Code Rewriting

### v1

@@ Abstract.
Large Language Models (LLMs) have exhibited remarkable proficiency in generating code. However, the misuse of LLM-generated (Synthetic) code has prompted concerns within both educational and industrial domains, highlighting the imperative need for the development of synthetic code detectors. Existing methods for detecting LLM-generated content are primarily tailored for general text and often struggle with code content due to the distinct grammatical structure of programming languages and massive ”low-entropy” tokens. Building upon this, our work proposes a novel zero-shot synthetic code detector based on the similarity between the code and its rewritten variants. Our method relies on the intuition that the differences between the LLM-rewritten and original codes tend to be smaller when the original code is synthetic. We utilize self-supervised contrastive learning to train a code similarity model and assess our approach on two synthetic code detection benchmarks. Our results demonstrate a notable enhancement over existing synthetic content detectors designed for general texts, with an improvement of 20.5% in the APPS benchmark and 29.1% in the MBPP benchmark.
@@ 1. Introduction
Large Language models (LLM) pretrained on code have shown remarkable capability in understanding and generating code (Chen et al., 2021; Fried et al., 2022; Li et al., 2023; Nijkamp et al., 2022). Those Code LLMs can function as professional coding assistants for programmers, offering intelligent code completion and document generation capabilities, such as Github Copilot. Furthermore, recent advancements in instruction tuning and AI alignment research (Wei et al., 2022a; Ouyang et al., 2022) have facilitated the development of general-purpose conversational LLMs (Chowdhery et al., 2022; OpenAi, 2022; OpenAI, 2023; Touvron et al., 2023). GPT-4, one of the most impressive conversational LLM released by OpenAI, can provide high-quality responses to general human requests in a zero-shot manner, including requests for generating code implementation for detailed coding specifications (Bubeck et al., 2023).
@@
The breakthrough of LLM has dramatically improved the efficiency of the coding process and significantly lowered the barrier to entry for programming. However, it also raises concerns about the misuse of LLM-generated code. In programming education, students have already used LLMs to write solutions for coding assignments and exams (Kazemitabaar et al., 2023; Denny et al., 2024). A recent study has indicated that GPT-4 can achieve the average human level performance in solving LeetCode problems (Bubeck et al., 2023), which makes cheating with the use of LLMs in programming exams very attractive for students, and also makes it difficult to assess students’ programming abilities in education. Besides educational domain, in real industrial applications, there is a high demand for code security, and LLM-generated code often contains security vulnerabilities (He and Vechev, 2023). An evaluation (Pearce et al., 2022) discovered that, in various security-relevant scenarios, 40% of Copilot-generated programs contain dangerous vulnerabilities. Therefore, if the code is LLM-generated, it requires even more rigorous review.
@@
Figure 1. Token Entropy Distribution of Natural Texts and Code. We use 500 documents from XSum (Narayan et al., 2018) dataset for token entropy estimation and 770 human-written code for code token entropy estimation.
@@
Building upon the aforementioned critical real-world scenarios, developing a synthetic code detector is curcial for addressing concerns regarding the misuse of LLM-generated code. Although there have been some efforts to detect LLM-generated pure text (Gehrmann et al., 2019; Mitchell et al., 2023; Zellers et al., 2019; Ippolito et al., 2020; Zhong et al., 2020; Pu et al., 2023), none of them have specifically focused on code content. The question of whether existing detection methods for general texts remain effective when applied to code content is still unknown. However, unfortunately, the answer to this question is negative. Even state-of-the-art detection methods, such as GLTR (Gehrmann et al., 2019) and DetectGPT (Mitchell et al., 2023), exhibit a significant decline in detection performance when applied to the code domain, as revealed by our experiments. Our findings also align with a recent evaluation study of AIGC detectors on code content (Wang et al., 2023).
@@
Figure 2. Large Language Models can only make minor modifications to the synthetic code. We use the StarChat model (Li et al., 2023) to rewrite the original code and observe that the similarity between the rewritten and original code is much higher when the original code is sampled from GPT-3.5-Turbo.
@@
Through our meticulous experiments, we identify the primary reason lies in the core logic of existing state-of-the-art methods for text detection, which is based on the log probability of tokens. Large Languages models assign higher log probability to the tokens they generate because these tokens are considered by the model to the most confident in its self-generated content. Consequently, existing detection methods leverage this characteristic to differentiate between model-generated text and human-written text, achieving high detection accuracy. However, in the domain of code, this approach exposes its inherent weaknesses. The main reason is that, compared to natural language, code exhibits a more uniform grammatical structure. In specific programming language and the given context, the subsequent code token is often deterministic or has limited options. For example, in the Python language, functions always start with ”def”, and a ”:” must follow ”def funciton_name()”. We refer to these tokens as ”low-entropy” tokens. However, there exists a large number of such ”low-entropy” tokens in code, reducing the actual effective tokens for SOTA text detection methods to differentiate between human-written code and model-generated code, resulting a statistically insignificant differences in distinguishing between these two types of code. We investigate the code token extropy distribution and the token entropy distribution (both estimated by Llama-13B) in Figure 1. The distrubution of text token entropy is more dispersed than code, and more than 60% of code tokens have entropy smaller than 1.
@@
Therefore, to address the detection challenge between model-generated code and human-written code in the code domain, it is essential to abandon the current token log probabilities perspective. Adopting a more holistic, global view, on the other hand, opens up our thinking. Through an essential observation, as illustrated in Figure 2, that Large language Model, when tasked with rewriting model-generated code produced by either themselves or other LLMs, tend to exhibit remarkably similar coding patterns in their rewritten output compared to the original code. Conversely, when tasked with rewriting human-written code, the LLM-rewritten code diverges more significantly from the original code. In Figure 2, in the top left corner is the original code (model-generated), which, after beging rewritten by LLMs, yields the code in the top right corner. It’s evident that the two are strikingly similar, with matching portions highlighted in yellow. Conversely, the bottom left corner showcases a human-written code. to the same LLM rewriting process, we obtain the code in the bottom right corner, where we observe a substantial number of differences.
@@
This observation inspired our hypothesis: the similarity between the original and rewritten code can be used as an indicator for detecting synthetic code. Furthermore, the generation of rewritten code and the utilization of similarity as an indicator allows detection methods to be independent of likelihood estimation, thereby addressing the root cause of failure in inaccurate likelihood estimation for code content. Based on this, we propose a zero-shot model-generated code detector using Code Rewriting and similarity Measurement.
@@
Our approach comprises two crucial steps: Code Rewriting and Similarity Measurement. When confronted with a piece of code, whether model-generated or human-written, we first employ a large language model to rewrite it using an appropriate prompt template. This process yields both the original code and its corresponding rewritten version. Subsequently, to accurately gauge the similarity between the original and rewritten code, we develop a code similarity model that predicts a similarity score for the given code pair. Finally, employing the cosine similarity metric, we estimate the expected similarity score by sampling m rewritten codes following the identical procedure and averaging their scores.
@@
We constructed two model-generated code detection benchmarks utilizing APPS (Hendrycks et al., 2021) and MBPP (Austin et al., 2021), representing use cases in the framework of synthetic code exploitation. To confirm that our intuition and hypothesis are universally applicable to all types of code generation tools, we generate synthetic code using two open-sourced LLMs (CodeLlama (Rozière et al., 2023), StarChat (Li et al., 2023)) and two proprietary code generation APIs (GPT-3.5, GPT-4) for detection.The experimental results demonstrate that the methods proposed in our study notably outperform SOTA benchmarks. Specially, our approach achieve 20.5% improvement in detection performance on the APPS dataset and 29.1% improvement in MBPP dataset. Furthermore, from a practicality and minimal resource requirement standpoint, our method only necessitates the capability to perform inference with or access to their APIs, in contrast to the previous text detection methods that require knowledge of token log probabilities.
@@
To facilitate further research, we have made our dataset, code and trained code similarity model checkpoint publicly available. Overall, the main contributions of this paper can be outlined as follows:
@@
  • We have identified a performance gap when applying synthetic content detectors designed for general texts to code content. Following a detailed analysis, we have uncovered the reasons for this discrepancy.
  • We proposed a zero-shot synthetic content detection tool specially designed for code content, applicable to both open-sourced code Large Language Models and closed-source LLMs like ChatGPT/GPT-4, which solely offer APIs.
@@
  • We proposed a novel detection approach utilizing code rewriting and similarity measurement, which effectively addressed the unique challenges for synthetic code detection.
@@
  • We conducted extensive experiments and analysis to demonstrate the effectiveness of our approach, showing significant improvements in accuracy compared to existing methods.
@@ 2.1. LLMs for Code Generation
With the rapid rise in popularity of large language models, an increasing number of studies have begun focusing on code large language models aimed at achieving automatic software engineering. Those code large language models are continuely pre-trained on large-scale code corpus from an initial pre-trained model on text with auto denoising (Wang et al., 2021a) or causal language modeling task (Chen et al., 2021; Li et al., 2023; Nijkamp et al., 2022; Fried et al., 2022). A crucial capability of large code models is code generation. Code generation is a left-to-right decoding process that utilizes functional requirements or code context as a prompt. Pioneering works such as CodeX (Chen et al., 2021), AlphaCode (Li et al., 2022a), CodeGeex (Zheng et al., 2023), SantaCoder (Allal et al., 2023), and GPT-4 (Achiam et al., 2023) have all demonstrated powerful code generation capabilities. Recent advances in general purpose LLMs (Chowdhery et al., 2022; OpenAi, 2022; OpenAI, 2023; Touvron et al., 2023) further enhanced their understanding of human instructions, resulting in the generation of code that better meets the requirements and exhibits higher quality.
As code large language models continue to mature, more and more people from various industries are integrating these models into their daily lives. According to a community survey led by StackOverflow (2023), 44% of experienced developers and 55% of beginners already used AI coding assistant, and most of them use two tools, ChatGPT (83%) and Github Copilot (56%). Therefore, in certain areas such as education and industrial settings, automated detection of whether code is written by humans or generated by models has become increasingly important.
@@ 2.2. Detection of Synthetic Text
Detecting AI-generated (synthetic) text has been studied before the emergency of LLM. The main works on synthetic text detection have two lines. One research line formulates detection as a binary classification problem by collecting synthetic texts from generative models and training a supervised model based on pre-trained transformers (Zellers et al., 2019; Ippolito et al., 2020; Zhong et al., 2020) or other neural models (Bakhtin et al., 2019; Uchendu et al., 2020). Another research line detects synthetic text by designing zero-shot metrics. Those metrics measure the relationship between a given text and the text distribution of generative models. Gehrmann et al. (Gehrmann et al., 2019) claims that the synthetic texts are sampled from the head of generative models’ distribution, so the average log probability score under the generative model $\mathcal{G}$ of a given text, i.e., $\frac{1}{L}\sum_{i=1}^{L}log(P_{\mathcal{G}}(x_{i}|x_{1},...,x_{i-1}))$, can be a simple and effective zero-shot metric (called GLTR). Su et al. (2023) proposed two metric, LRR and NPR, where the former combines log-rand and log-likelihood to better magnify the differences between human-written and machine-written text, while the later is primarily based the idea that the log rank of machine-generated texts should be more sensitive to smaller perturbations. Mitchell et al. (2023) further improved the GLTR metrics by proposing that the synthetic texts lie at the probability curvature, i.e., the local maximum of the generative model’s distribution. They use the average token probability disparity between the given text and perturbed texts to detect whether the given text is located at the probability curvature of the generative model. DetectGPT4Code (Yang et al., 2023) introduces three key modifications to the original DetectGPT for adaptation to the code domain. Overall, all of these methods converge on the same idea: large language models tend to generate tokens with higher confidence. Therefore, effective differentiation can be achieved by statistically analyzing token log probabilities.
In addition to detection methods, some approaches have explored adding watermarks to the generated text during the generation process behind LLM services(Kirchenbauer et al., 2023; Yoo et al., 2023). Through watermark extraction, it becomes feasible to discern whether a given text originates from a model. However, it’s worth noting that the inclusion of a watermark algorithm could potentially compromise the quality of the generated text. Consequently, it cannot be presumed that all users or service providers will be inclined to produce watermarked text.
3. Method
@@ 3.1. Problem Definition
We focus on the zero-shot synthetic code detection problem. The synthetic code detection problem involves determining whether a code snippet $x$ is generated from a generative model or a API service (such as ChatGPT). The term ”zero-shot” refers to the scenario that we do not have access to a dataset with labeled synthetic code and human-written code for model training. All the previous zero-shot synthetic text detection methods, including GLTR and DetectGPT consider the ”white-box” setting, where log probability score of generative model is accessible. In the synthetic code detection problem, we argue that the ”white-box” assumption is too strong because in many cases, some commercial models only provide the generated content and not provide the log probability score. Therefore, we do not assume access to the log probability score and study the ”black-box” setting, where only generated content is available.

### v2

@@ Abstract.
Large Language Models (LLMs) have exhibited remarkable proficiency in generating code. However, the misuse of LLM-generated (synthetic) code has prompted concerns within both educational and industrial domains, highlighting the imperative need for the development of synthetic code detectors. Existing methods for detecting LLM-generated content are primarily tailored for general text and often struggle with code content due to the distinct grammatical structure of programming languages and massive ”low-entropy” tokens. Building upon this, our work proposes a novel zero-shot synthetic code detector based on the similarity between the code and its rewritten variants. Our method relies on the intuition that the differences between the LLM-rewritten and original codes tend to be smaller when the original code is synthetic. We utilize self-supervised contrastive learning to train a code similarity model and assess our approach on two synthetic code detection benchmarks. Our results demonstrate a significant improvement over current synthetic content detectors designed for general texts, achieving a 20.5% increase in AUROC on the APPS benchmark and a 29.1% increase on the MBPP benchmark.
@@ 1. Introduction
Large Language models (LLMs) pre-trained on code have shown remarkable capability in understanding and generating code (Chen et al., 2021; Fried et al., 2022; Li et al., 2023; Nijkamp et al., 2022). Those code LLMs can function as professional coding assistants for programmers, offering intelligent code completion and document generation capabilities, such as Github Copilot [^1] https://copilot.github.com. Furthermore, recent advancements in instruction tuning and AI alignment research (Wei et al., 2022a; Ouyang et al., 2022) have facilitated the development of general-purpose conversational LLMs (Chowdhery et al., 2022; OpenAi, 2022; OpenAI, 2023; Touvron et al., 2023). GPT-4, one of the most impressive conversational LLM released by OpenAI, can provide high-quality responses to general human requests in a zero-shot manner, including requests for generating code implementation for detailed coding specifications (Bubeck et al., 2023).
@@
The breakthrough of LLM has dramatically improved the efficiency of the coding process and significantly lowered the barrier to entry for programming. However, it also raises concerns about the misuse of LLM-generated code. In programming education, students have already used LLMs to write solutions for coding assignments and exams (Kazemitabaar et al., 2023; Denny et al., 2024). A recent study has shown that GPT-4 can achieve the average human-level performance in solving LeetCode problems (Bubeck et al., 2023). This makes cheating with the use of LLMs in programming exams very attractive for students and also makes it difficult to assess students’ programming abilities in education. Besides the educational domain, there is a significant demand for code security in real industrial applications, and LLM-generated code often contains security vulnerabilities (He and Vechev, 2023). An evaluation (Pearce et al., 2022) revealed that in various security-relevant scenarios, 40% of Copilot-generated programs contain dangerous vulnerabilities. Therefore, if the code is synthetic, it requires an even more rigorous review.
@@
Figure 1. Token Entropy Distribution of Natural Texts and Code. We utilize 500 documents from XSum (Narayan et al., 2018) dataset for token entropy estimation and 770 human-written code samples for code token entropy estimation.
@@
Building upon the aforementioned critical real-world scenarios, developing a synthetic code detector is crucial for addressing concerns regarding the misuse of LLM-generated code. Although there have been some efforts to detect LLM-generated pure text (Gehrmann et al., 2019; Mitchell et al., 2023; Zellers et al., 2019; Ippolito et al., 2020; Zhong et al., 2020; Pu et al., 2023), none of them have specifically focused on code content. The effectiveness of existing detection methods for general texts when applied to code content remains unknown. Unfortunately, our experiments, detailed in Section 5.1, reveal that even state-of-the-art detection methods, such as GLTR (Gehrmann et al., 2019) and DetectGPT (Mitchell et al., 2023), exhibit a significant decline (approximately 32%) in detection performance when applied to the code domain.
@@
Figure 2. Large Language Models can only make minor modifications to the synthetic code. We use the StarChat model (Li et al., 2023) to rewrite the original code and observe that the similarity between the rewritten and original code is much higher when the original code is generated from GPT-3.5-Turbo.
@@
To analyze the reasons for the failure, our meticulous experiments reveal that the primary issue lies in the core logic of existing state-of-the-art methods for text detection, which rely on the statistical log probability of tokens. LLMs assign higher log probabilities to the tokens they generate due to their high confidence in these self-generated tokens. Consequently, existing detection methods leverage this characteristic to differentiate between model-generated text and human-written text, achieving high detection accuracy. However, in the domain of code, this approach exposes its inherent weaknesses. The main reason is that, compared to natural language, code exhibits a more uniform grammatical structure. In specific programming languages and the given code context, the subsequent code token is often deterministic or has limited options. For example, in Python, functions always start with ”def”, and a ”:” must follow the function signature ”def funciton_name()”. We refer to these tokens as ”low-entropy” tokens. However, there exists a large number of such ”low-entropy” tokens in code domain, greatly reducing the actual effective tokens for SOTA text detection methods to differentiate between human-written code and model-generated code, resulting in statistically insignificant differences between the two. We investigate the code token entropy distribution and the token entropy distribution (both estimated by Llama-13B) in Figure 1. The distribution of text token entropy is more dispersed than code, and more than 70% of code tokens have entropy smaller than 1.
@@
Therefore, to address the detection challenge between human-written code and synthetic code, it is essential to abandon the current token log probabilities perspective. Adopting a more holistic, global view, on the other hand, opens up our thinking. Through a global and essential observation, we found that large language model, when tasked with rewriting synthetic code produced by either themselves or other LLMs, tend to exhibit remarkably similar coding patterns in their rewritten output compared to the original code. Conversely, when tasked with rewriting human-written code, the LLM-rewritten code diverges more significantly from the original code. As shown in Figure 2, the top left corner shows the original code (synthetic), which, after being rewritten by LLMs, yields the code in the top right corner. It’s evident that the two are strikingly similar, with matching portions highlighted in orange. Conversely, the bottom left corner showcases human-written code. After undergoing the same LLM rewriting process, the LLM rewriting code in the bottom right corner shows a substantial number of differences compared to the original human-written code.
@@
This observation inspired our hypothesis: the similarity between the original and rewritten code can be used as an indicator for detecting synthetic code. Furthermore, the generation of rewritten code and the utilization of similarity as an indicator allows detection methods to be independent of statistically token likelihood estimation, thereby addressing the root cause of failure in current text detector for the code context. Based on this hypothesis, we propose a zero-shot synthetic code detector using Code Rewriting and Similarity Measurement.
@@
Our method comprises two crucial steps: Code Rewriting and Similarity Measurement. When confronted with a piece of code, whether synthetic or human-written, we first employ a large language model to rewrite it using an appropriate prompt template. This process yields a pair consisting of the original code and its corresponding rewritten version. To accurately gauge the similarity between the original and rewritten code, we then develop a code similarity model that predicts a similarity score for the given code pair. Finally, using the cosine similarity metric, we estimate the expected similarity score by sampling $m$ rewritten codes following the identical procedure and averaging their scores.
@@
Due to the lack of synthetic code detection datasets, we construct two synthetic code detection datasets utilizing APPS (Hendrycks et al., 2021) and MBPP (Austin et al., 2021), representing use cases in the framework of synthetic code exploitation. To confirm that our hypothesis and method are universally applicable to all types of code generation tools, we generate synthetic code using two open-sourced LLMs (CodeLlama (Rozière et al., 2023), StarChat (Li et al., 2023)) and two proprietary code generation APIs (GPT-3.5 (OpenAi, 2022), GPT-4 (OpenAI, 2023)) for detection. The experimental results demonstrate that our proposed method notably outperforms current SOTA methods. Specifically, our approach achieves a 20.5% improvement in detection performance (AUROC) on the APPS dataset and a 29.1% improvement on the MBPP dataset. Furthermore, from a practicality and minimal resource/permission requirement standpoint, our method only necessitates the capability to perform LLM inference or access to their APIs. This is in contrast to the previous detection methods that require knowledge of token log probabilities.
@@
To facilitate further research, we have made our dataset, code, and trained code similarity model checkpoint publicly available [^2] https://anonymous.4open.science/r/code-detection-6B35. Overall, the main contributions can be outlined as follows:
@@
  • We identify a performance gap when applying synthetic content detectors designed for general texts to the code domain. Through detailed analysis, we uncovered the reasons for this discrepancy.
@@
  • We propose a novel detection approach that utilizing code rewriting and similarity measurement, effectively addressing the unique challenges of synthetic code detection.
  • Our proposed zero-shot synthetic code detector is applicable to both open-sourced code LLMs and closed-source LLMs like ChatGPT/GPT-4, which solely offer APIs.
@@
  • Extensive experiments and analysis demonstrate the effectiveness of our method, revealing significant improvements in accuracy and robustness compared to existing methods.
@@ 2.1. LLMs for Code Generation
With the rapid rise in the popularity of LLMs, an increasing number of studies have begun focusing on Code LLMs aimed at achieving automatic software engineering. Those Code LLMs are continually pre-trained on large-scale code corpus from an initial pre-trained model on text with auto-denoising (Wang et al., 2021a) or causal language modeling task (Chen et al., 2021; Li et al., 2023; Nijkamp et al., 2022; Fried et al., 2022). A crucial capability of Code LLMs is code generation. Code generation is a left-to-right decoding process that utilizes functional requirements or code context as a prompt. Pioneering works such as CodeX (Chen et al., 2021), AlphaCode (Li et al., 2022a), CodeGeex (Zheng et al., 2023), SantaCoder (Allal et al., 2023), and GPT-4 (OpenAI et al., 2024) have all demonstrated powerful code generation capabilities. Recent advances in general purpose LLMs (Chowdhery et al., 2022; OpenAi, 2022; OpenAI, 2023; Touvron et al., 2023) further enhanced their understanding of human instructions, resulting in the generation of code that better meets the requirements and exhibits higher quality.
As code large language models continue to mature, more and more people from various industries are integrating these models into their daily lives. According to a community survey led by StackOverflow (2023), 44% of experienced developers and 55% of beginners already used AI coding assistant and most of them use two tools, ChatGPT (83%) and Github Copilot (56%). Therefore, in certain areas such as education and industrial settings, automated detection of whether code is written by humans or generated by models has become increasingly important.
@@ 2.2. Detection of Synthetic Text
Detecting AI-generated (synthetic) text has been studied before the emergency of LLM. The main works on synthetic text detection follow two lines of research. One research line formulates detection as a binary classification problem by collecting synthetic texts from generative models and training a supervised model based on pre-trained transformers (Zellers et al., 2019; Ippolito et al., 2020; Zhong et al., 2020) or other neural models (Bakhtin et al., 2019; Uchendu et al., 2020). Another research line detects synthetic text by designing zero-shot metrics. These metrics measure the relationship between a given text and the text distribution of generative models. Gehrmann et al. (Gehrmann et al., 2019) claim that synthetic texts are sampled from the head of generative models’ distribution, so the average log probability score under the generative model $\mathcal{G}$ of a given text, i.e., $\frac{1}{L}\sum_{i=1}^{L}log(P_{\mathcal{G}}(x_{i}|x_{1},...,x_{i-1}))$, can be a simple and effective zero-shot metric (called GLTR). Su et al. (2023) proposed two metrics, LRR and NPR, where the former combines log-rand and log-likelihood to better magnify the differences between human-written and machine-generated text, while the latter is primarily based on the idea that the log-rank of machine-generated texts should be more sensitive to smaller perturbations. Mitchell et al. (2023) further improved the GLTR metrics by proposing that synthetic texts tend to occupy negative curvature regions of the model’s log probability function, i.e., the local maximum of the generative model’s distribution. They use the average token probability disparity between the given text and perturbed texts to detect whether the given text is located at the negative curvature regions of $log\ p(x)$. Overall, all of these methods converge on the same idea: LLMs tend to generate tokens with higher confidence. For tokens generated by LLMs themselves, there will be assigned a higher log probability. Therefore, effective differentiation can be achieved by statistically analyzing token log probabilities.
In addition to detection methods, some approaches have explored adding watermarks to the generated text during the generation process behind LLM services(Kirchenbauer et al., 2023; Yoo et al., 2023). Through watermark extraction, it becomes feasible to discern whether a given text originates from a model. However, it’s worth noting that the inclusion of a watermarking algorithm could potentially compromise the quality of the generated text. Consequently, it cannot be presumed that all users or service providers will be inclined to produce watermarked text.
3. Methodology
@@ 3.1. Problem Definition
We focus on the zero-shot synthetic code detection problem. The synthetic code detection problem involves determining whether a code snippet $x$ is generated from a generative LLM or an API service (such as ChatGPT). The term ”zero-shot” refers to the scenario where we do not have access to a dataset with labeled synthetic code and human-written code for model training. Notably, all previous zero-shot synthetic text detection methods, including GLTR and DetectGPT, consider the ”white-box” setting, where log probability score of the generative model is accessible. However, in the synthetic code detection problem, we argue that the ”white-box” assumption is too strong because in many cases, some commercial models only provide the generated content and do not provide the log probability score. Therefore, we do not assume access to the log probability score and study the ”black-box” setting, where only generated content is available.